# A joint optimization approach to identifying sparse dynamics using least squares kernel collocation

### v1

@@ 5 Discussion and Conclusion
The least squares relaxation has been applied in the PDE inverse problem setting [tristan_pde], where it was found to mitigate the effects of local minima [vanLeeuwen2013Mitigating]. It was further analyzed in [tristan_analysis], which presented a representer theorem for relaxations of parametric linear PDEs and cast the inversion problem as minimizing a special RKHS norm. The least squares approach is also common in the ML literature, most notably through physics informed neural networks [Raissi2019-so], as well as approaches to learning differential equations and parameter estimation [sun2022bayesiansplinelearningequation, Chen2021, ijcai2021p283, Fung2025-odr, jalalian2025dataefficientkernelmethodslearning, rudy_soft]. In our case, the additional error associated to the ODE from applying a least-squares approach can be beneficial to the overall learning process, rather than a limitation. We prefer the relaxation approach to alternatives, such as reduced or shooting methods that view the ODE as a constraint and enforce feasibility from the start, which can be seen in neural ODEs, inverse problems, and optimal control [kidger2021on, Chen2018, Givoli2021-bc, Lions1971], or formulations that enforce the hard constraint within an all-at-once formulation such as [Hokanson2023-sr], which optimizes by solving the discretized KKT conditions.
@@
The RKHS regularized least squares approach that we take has two major advantages over reduced approaches. First, state estimates for any length of time are well defined for a much broader class of ODEs, which can be seen from the existence of bounded solutions in Theorem 2.2. Further, the state estimates are allowed to be self-correcting in that the additional data misfit terms may inform the state estimates. In contrast, in reduced approaches, certain parameter values may lead to finite time divergence of the ODE. Second, this approach improves the regularity of the objective with respect to the system parameters $\boldsymbol{\theta}$. For chaotic dynamics, we can generically expect that small perturbations to parameters of the system will lead to exponentially large changes over long periods of time relative to the Lyapunov exponent. Several works have observed that this leads to extremely ill-conditioned and non-convex objectives in reduced formulations, especially in chaotic regimes where the mismatch between simulated and observed trajectories can grow exponentially with time [ribeiro2020smoothness, Chakraborty2024multistep, Carbonell2015multishooting]. As a result, gradient-based optimization often stalls, converges to spurious local minima, or fails entirely. These issues can be mitigated significantly through multiple-shooting and multipoint-penalty methods [Chakraborty2024multistep, Carbonell2015multishooting, Chung2022-hz], which break a long trajectory into shorter segments, while simulating within each segment. The full space approaches considered in this work can, in a sense, be seen as the limit of many segments in a multiple-shooting method, but we leave a more systematic comparison to future work.
@@
The least squares relaxation avoids simulating forward trajectories altogether and instead distributes the residual over time, substantially smoothing the optimization landscape and improving robustness. Further, it makes the objective a linear least squares regression problem in the coefficients $\boldsymbol{\theta}$ for any fixed state estimate $\boldsymbol{x}$. This natural perspective also appears in [tristan_analysis] in the context of recovering a coefficient field in a PDE-constrained inverse problem. [Fung2025-odr] points out the connection to orthogonal distance regression in all-at-once approaches, casting optimization with respect to the state variables as adjusting the inputs to correct an error-in-variables bias. Indeed, in the linear case, such an approach appears as early as 1950 in Householder’s method for exponential fitting [householder1950prony], where the error-in-variables issue in Prony’s method is corrected by jointly fitting state estimates and discrete time dynamics.
@@
While one can control the error in terms of the size of the residual $\boldsymbol{r}$ and a Lipschitz assumption on $f$ via classical error analysis for ODEs using Grönwall’s inequality, this will generally result in overly pessimistic error estimates which grow exponentially in time as it assumes worst case behavior of $\boldsymbol{r}$. Rather, $\boldsymbol{r}$ largely pushes the trajectory to match the data, so its effects should be small when the model is well specified, and large in misspecified cases, where it is necessary. By assimilating the observation data over a long trajectory, the error can be compared to that of Kalman filtering/smoothing (see e.g. [aravkin2017generalized]) under very low system noise and a reasonably accurate system model.
In our case, the regularity of $\boldsymbol{r}$ follows from the smoothness of $\boldsymbol{x}$ lent by the RKHS regularization and continuity of $f$, noting that $\boldsymbol{r}(t)=\frac{d^{p}}{dt^{p}}\boldsymbol{x}(t)-f(D[\boldsymbol{x}](t);\boldsymbol{\theta})$, the difference of two continuous functions. However, without regularization of $\boldsymbol{x}$, there is no reason to believe that $\boldsymbol{r}$ is continuous, so $\boldsymbol{x}$ will only be piecewise $p$-times differentiable; for first order ODEs, one should expect kinks in the state estimates without higher order regularization. The closely related approach of ODR-BINDy [Fung2025-odr] interprets the least squares approach as a stochastic relaxation, modeling error in the ODE as independent white noise. However, it can be difficult to rigorously interpret the noise stochastically taking the function space limit of inferring $\boldsymbol{x}(\cdot)$; in the random setting and without regularization, the covariance of the residuals becomes increasingly important as the mesh is refined and $\boldsymbol{r}(t)$ is considered as a random process. In the case that the residuals are uncorrelated (e.g. $\boldsymbol{r}(t),\boldsymbol{r}(t^{\prime})$ are independent for $t\neq t^{\prime}$ ), then either the pointwise variances have to blow up, or the noise completely washes out under the integral. The probabilistic model can be seen as applying maximum a posteriori estimation to the stochastic differential equation
@@
In this model, $\frac{d}{dt}\boldsymbol{x}$ will not be defined pointwise, possessing only the regularity of white noise. Of course, these considerations do not manifest computationally; the error terms are end up correlated, there is no “real” Brownian motion happening, and most reasonable discretizations will implicitly regularize the trajectory estimate. However, this does raise the question of determining a computationally effective and complete stochastic model in the vein of probabilistic numerics [probnum], as well as the idea of penalizing the residual in a stronger norm than $L^{2}$.
Conclusion
We proposed and demonstrated the effectiveness of JSINDy in both system identification and filtering by performing both jointly. We addressed a number of challenging problems in continuous time system within a single framework based on collocation in RKHSs, including scarse sampling, far below what has been considered in previous works, partially observed states, and higher order dynamics. Building on other recent advancements in all-at-once system identification methods [Fung2025-odr, Chen2021, ijcai2021p283, Ribera2022-nz, Hokanson2023-sr], this work provides more strong evidence for the superiority of all-at-once methods for system identification, and demonstrates significant flexibility. In short, the assumption that the data satisfy some dynamics is a very good prior to use when it is true. Overall, the RKHS formulation provides a flexible and effective tool for modeling while providing a principled approach to regularization, resulting in stable optimization, and allowing for seamlessly incorporating prior assumptions, data, and dynamics.
Currently, the primary drawback of JSINDy is that dense kernel matrices arise from the global approximation, leading to cubic complexity in the trajectory length. This can, however, be mitigated, either through a state-space representation of Matèrn Gaussian processes as in [Saa12], which essentially amounts to a finite difference approximation to the associated Sobolev norm, and would result in banded Jacobian matrices.
Further improvements are available on the side of optimization. The objective function in eq. 8 is amenable to partial minimization [Aravkin2024-bz, O-Leary2013-mg, Golub1973-el] with respect to the coefficients $\boldsymbol{\theta}$ in solving the fixed-support nonlinear least squares problems. Our approach of alternating with an algorithmic sparsifier another choice which could use more investigation. In particular, incorporating Bayesian evidence as a model selection criteria could be highly effective, while raising interesting questions for efficiently solving discrete optimization problems that arise [Fung2025-odr, Fung2025-aa, Klishin2025-pl].
@@ Acknowledgments
AH and BH acknowledge support from the National Science Foundation under awards 2208535 (Machine Learning for Bayesian Inverse Problems) and 2337678 (CAREER: Gaussian Processes for Scientific Machine Learning: Theoretical Analysis and Computational Algorithms). AH acknowledges support from a Carl E. Pearson Fellowship. JNK was supported in part by the NSF AI Institute for Dynamical Systems (dynamicsai.org), grant 2112085. The numerical experiments involving ODR-BINDy were completed on Hyak, UW’s high performance computing cluster, which is funded by the UW student technology fee.
[…]
@@ A.1 RKHS and representer theorems
We first give a version of the representer theorem in the Hilbert space setting which is appropriate for Theorem 2.1. These theorems are well known and appear in various forms throughout the literature;we state it for completeness. The variant that we present can be found in [unifying-representer, Owhadi2019-xh, Scholkopf2001-co], and similar results appear in [CHEN2021110668, ZhouDerivative, kanagawa2018gaussianprocesseskernelmethods].
Theorem A.1.
@@
where $J:\mathbb{R}^{+}\mapsto\mathbb{R}\cup\{\infty\}$ is strictly increasing, $w_{i}$ are bounded linear functionals, and $f:\mathbb{R}^{m}\mapsto\mathbb{R}\cup\{\infty\}$. Then if $F$ admits a minimizer $u^{*}$ with $F(u^{*})<\infty$, then
@@ A.2 Matérn kernels and Sobolev spaces
Here, we review some prerequisite RKHS and Sobolev space theory, giving more details on the kernel functions we use, and discussing the smoothness of functions in RKHS $\mathcal{X}$ associated to the kernels we define in eq. 9. Much of this will pull directly from [kanagawa2018gaussianprocesseskernelmethods, Wendland2010-gl] where we direct the interested reader for additional details. First, recall from eq. 9 that we take the sum of a constant kernel with a Matérn kernel
@@
From [kanagawa2018gaussianprocesseskernelmethods, Ex. 2.6], the RKHS $\mathcal{H}_{k}$ associated to this kernel is norm equivalent to the Sobolev space $H^{\nu+\frac{1}{2}}([0,T])$, functions which are $\nu+\frac{1}{2}$ times differentiable in the mean-square sense. Compactness of the domain means that adding a constant to the kernel function does not affect the equivalence. By the Sobolev embedding [evansPDE, Thm. 5.6.6, p. 270], this implies that $\mathcal{H}_{k}$ is also compactly embedded in the Hölder space $C^{p}([0,T])$ for $p<\nu$.
When $\nu$ is a half integer (i.e. $\nu=n+\frac{1}{2}$), $h_{\nu}$ reduces to the product of a polynomial and exponential in $r$.
@@
Stably computing high order derivatives of this $h_{\nu}$ can still be challenging due to the absolute value inside of $h_{\nu}$, and directly applying autodiff to this representation can cause catastrophic cancellation. However, for $j<2\nu$, $\frac{d^{j}}{dt^{j}}h_{\nu}(\frac{|t-t^{\prime}|}{\ell})$, can also be written as the product of a polynomial and exponential function; we compute this representation of the derivatives in SymPy[sympy], and then convert it Jax using Sympy2Jax [Kidger_sympy2jax_2025].
@@ A.3 Cholesky factorization and choice of basis
We perform a linear change of variables in order to simplify bookkeeping and mitigate some of the effects of ill-conditioning and numerical roundoff on convergence by reparametrizing with the Cholesky factors of $\boldsymbol{K}$. That is, write
@@
It is clear that these are interchangable mathematically. Computationally, they relegate some of the bookkeeping with tracking the kernel matrix to an initial Cholesky factorization, which, in our implementation, is absorbed into the parametrization, though for simplicity here, we continue the exposition using the original variables. The fact that this mitigates some of the effects of ill-conditioning was somewhat mysterious to us, as it still entails computing Cholesky factorizations of ill-conditioned matrices.
@@ Appendix B Computational tools
We perform all of our computations using Jax [jax2018github], and solve ODEs using the Tsit5 [tsitouras2011runge] solver with a PID controller for adaptive timestepping [hairer2002solving-ii, soderlind2002automatic] as recommended by, and implemented in Diffrax [kidger2021equinox, kidger2021on]. Comparisons with SINDy use the PySINDy package [Kaptanoglu2022]. Our experiments and current implementation are available in the GitHub repository [jsindyRepo]; an optimized implementation of JSINDy will soon be added to PySINDy. Our computations with JAX used an RTX-5000 Ada Generation GPU in a Lambda machine.

### v2

@@ 5 Discussion and conclusion
The least squares relaxation has been applied in the PDE inverse problem setting [53], where it was found to mitigate the effects of local minima [94]. It was further analyzed in [95], which presented a representer theorem for relaxations of parametric linear PDEs and cast the inversion problem as minimizing a special RKHS norm. The least squares approach is also common in the ML literature, most notably through physics-informed neural networks [72], as well as approaches to learning differential equations and parameter estimation [88, 22, 87, 28, 40, 77]. In our case, the additional error associated to the ODE from applying a least-squares approach can be beneficial to the overall learning process, rather than a limitation. We prefer the relaxation approach to alternatives, such as reduced or shooting methods that view the ODE as a constraint and enforce feasibility from the start, which can be seen in neural ODEs, inverse problems, and optimal control [48, 20, 32, 55], or formulations that enforce the hard constraint within an all-at-once formulation such as [38], which optimizes by solving the discretized KKT conditions.
@@
The RKHS regularized least squares approach that we take has two major advantages over reduced approaches. First, state estimates for any length of time are well defined for a much broader class of ODEs, which can be seen from the existence of bounded solutions in Theorem 2.2. Further, the state estimates are allowed to be self-correcting in that the additional data misfit terms may inform the state estimates. In contrast, in reduced approaches, certain parameter values may lead to finite time divergence of the ODE. Second, this approach improves the regularity of the objective with respect to the system parameters $\boldsymbol{\theta}$. For chaotic dynamics, we can generically expect that small perturbations to parameters of the system will lead to exponentially large changes over long periods of time relative to the Lyapunov exponent. Several works have observed that this leads to extremely ill-conditioned and non-convex objectives in reduced formulations, especially in chaotic regimes where the mismatch between simulated and observed trajectories can grow exponentially with time [75, 17, 16]. As a result, gradient-based optimization often stalls, converges to spurious local minima, or fails entirely. These issues can be mitigated significantly through multiple-shooting and multipoint-penalty methods [17, 16, 23], which break a long trajectory into shorter segments, while simulating within each segment. The full space approaches considered in this work can, in a sense, be seen as the limit of many segments in a multiple-shooting method, but we leave a more systematic comparison to future work.
@@
The least squares relaxation avoids simulating forward trajectories altogether and instead distributes the residual over time, substantially smoothing the optimization landscape and improving robustness. Further, it makes the objective a linear least squares regression problem in the coefficients $\boldsymbol{\theta}$ for any fixed state estimate $\boldsymbol{x}$. This natural perspective also appears in [95] in the context of recovering a coefficient field in a PDE-constrained inverse problem. [28] points out the connection to orthogonal distance regression in all-at-once approaches, casting optimization with respect to the state variables as adjusting the inputs to correct an error-in-variables bias. Indeed, in the linear case, such an approach appears as early as 1950 in Householder’s method for exponential fitting [39], where the error-in-variables issue in Prony’s method is corrected by jointly fitting state estimates and discrete time dynamics.
@@
While one can control the error in terms of the size of the residual $\boldsymbol{r}$ and a Lipschitz assumption on $f$ via classical error analysis for ODEs using Grönwall’s inequality, this will generally result in overly pessimistic error estimates which grow exponentially in time as it assumes worst case behavior of $\boldsymbol{r}$. Rather, $\boldsymbol{r}$ largely pushes the trajectory to match the data, so its effects should be small when the model is well specified, and large in misspecified cases, where it is necessary. By assimilating the observation data over a long trajectory, the error can be compared to that of Kalman filtering/smoothing (see e.g. [1]) under very low system noise and a reasonably accurate system model.
In our case, the regularity of $\boldsymbol{r}$ follows from the smoothness of $\boldsymbol{x}$ lent by the RKHS regularization and continuity of $f$, noting that $\boldsymbol{r}(t)=\frac{d^{p}}{dt^{p}}\boldsymbol{x}(t)-f(D[\boldsymbol{x}](t);\boldsymbol{\theta})$, the difference of two continuous functions. However, without regularization of $\boldsymbol{x}$, there is no reason to believe that $\boldsymbol{r}$ is continuous, so $\boldsymbol{x}$ will only be piecewise $p$-times differentiable; for first order ODEs, one should expect kinks in the state estimates without higher order regularization. The closely related approach of ODR-BINDy [28] interprets the least squares approach as a stochastic relaxation, modeling error in the ODE as independent white noise. However, it can be difficult to rigorously interpret the noise stochastically taking the function space limit of inferring $\boldsymbol{x}(\cdot)$; in the random setting and without regularization, the covariance of the residuals becomes increasingly important as the mesh is refined and $\boldsymbol{r}(t)$ is considered as a random process. In the case that the residuals are uncorrelated (e.g. $\boldsymbol{r}(t),\boldsymbol{r}(t^{\prime})$ are independent for $t\neq t^{\prime}$ ), then either the pointwise variances have to blow up, or the noise completely washes out under the integral. The probabilistic model can be seen as applying maximum a posteriori estimation to the stochastic differential equation
@@
In this model, $\frac{d}{dt}\boldsymbol{x}$ will not be defined pointwise, possessing only the regularity of white noise. Of course, these considerations do not manifest computationally; the error terms end up correlated, there is no “real” Brownian motion happening, and most reasonable discretizations will implicitly regularize the trajectory estimate. However, this does raise the question of determining a computationally effective and complete stochastic model in the vein of probabilistic numerics [36], as well as the idea of penalizing the residual in a stronger norm than $L^{2}$.
5.1 Conclusion
We proposed and demonstrated the effectiveness of JSINDy in both system identification and filtering by performing both jointly. We addressed a number of challenging problems in continuous time system identification within a single framework based on collocation in RKHSs, including scarce sampling, far below what has been considered in previous works, partially observed states, and higher order dynamics. Building on other recent advancements in all-at-once system identification methods [28, 22, 87, 76, 38], this work provides more strong evidence for the superiority of all-at-once methods for system identification, and demonstrates significant flexibility. In short, the assumption that the data satisfy some dynamics is a very good prior to use when it is true. Overall, the RKHS formulation provides a flexible and effective tool for modeling while providing a principled approach to regularization, resulting in stable optimization, and allowing for seamlessly incorporating prior assumptions, data, and dynamics.
Currently, the primary drawback of JSINDy is that dense kernel matrices arise from the global approximation, leading to cubic complexity in the trajectory length. This can, however, be mitigated through a state-space representation of Matérn Gaussian processes as in [79], which essentially amounts to a finite difference approximation to the associated Sobolev norm, and would result in banded Jacobian matrices.
Further improvements are available on the side of optimization. The objective function in eq. 8 is amenable to partial minimization [2, 67, 33] with respect to the coefficients $\boldsymbol{\theta}$ in solving the fixed-support nonlinear least squares problems. Our approach of alternating with an algorithmic sparsifier is another choice which could use more investigation. In particular, incorporating Bayesian evidence as a model selection criteria could be highly effective, while raising interesting questions for efficiently solving the discrete optimization problems that arise [28, 27, 50].
@@ Acknowledgments
AH and BH acknowledge support from the National Science Foundation under awards 2208535 (Machine Learning for Bayesian Inverse Problems) and 2337678 (CAREER: Gaussian Processes for Scientific Machine Learning: Theoretical Analysis and Computational Algorithms). AH acknowledges support from a Carl E. Pearson Fellowship. JMSH was partially supported by the Veterans Administration via the Post-9/11 GI Bill. JNK was supported in part by the NSF AI Institute for Dynamical Systems (dynamicsai.org), grant 2112085. The numerical experiments involving ODR-BINDy were completed on Hyak, UW’s high performance computing cluster, which is funded by the UW student technology fee.
@@ A.1 RKHS and representer theorems
We first give a version of the representer theorem in the Hilbert space setting which is appropriate for Theorem 2.1. These theorems are well known and appear in various forms throughout the literature; we state this for completeness. The variant that we present can be found in [4, 68, 83], and similar results appear in [21, 104, 44].
Theorem A.1 (Representer).
@@
where $J:\mathbb{R}^{+}\mapsto\mathbb{R}\cup\{\infty\}$ is strictly increasing, $\langle\cdot,w_{i}\rangle$ are bounded linear functionals, and $f:\mathbb{R}^{m}\mapsto\mathbb{R}\cup\{\infty\}$. Then if $F$ admits a minimizer $u^{*}$ with $F(u^{*})<\infty$, then
@@ A.2 Matérn kernels and Sobolev spaces
Here, we review some prerequisite RKHS and Sobolev space theory, giving more details on the kernel functions we use, and discussing the smoothness of functions in RKHS $\mathcal{X}$ associated to the kernels we define in eq. 9. Much of this will pull directly from [44, 98] where we direct the interested reader for additional details. First, recall from eq. 9 that we take the sum of a constant kernel with a Matérn kernel
@@
From [44, Ex. 2.6], the RKHS $\mathcal{H}_{k}$ associated to this kernel is norm equivalent to the Sobolev space $H^{\nu+\frac{1}{2}}([0,T])$, functions which are $\nu+\frac{1}{2}$ times differentiable in the mean-square sense. Compactness of the domain means that adding a constant to the kernel function does not affect the equivalence. By the Sobolev embedding thoerem [25, Thm. 5.6.6, p. 270], this implies that $\mathcal{H}_{k}$ is also compactly embedded in the Hölder space $C^{p}([0,T])$ for $p<\nu$.
When $\nu$ is a half integer (i.e. $\nu=p+\frac{1}{2}$), $h_{\nu}$ reduces to the product of a polynomial and exponential in $r$.
@@
Stably computing high order derivatives of this $h_{\nu}$ can still be challenging due to the absolute value inside of $h_{\nu}$, and directly applying autodiff to this representation can cause catastrophic cancellation. However, for $j<2\nu$, $\frac{d^{j}}{dt^{j}}h_{\nu}(\frac{|t-t^{\prime}|}{\ell})$, can also be written as the product of a polynomial and exponential function; we compute this representation of the derivatives in SymPy [63]. We then convert the resulting expressions to Jax [10] using Sympy2Jax [49].
@@ A.3 Cholesky factorization and choice of basis
In our numerical implementation, we perform a linear change of variables in order to simplify bookkeeping and mitigate some of the effects of ill-conditioning and numerical roundoff on convergence by reparametrizing with the Cholesky factors of $\boldsymbol{K}$. That is, write
@@
It is clear that these are interchangeable mathematically. Computationally, they relegate some of the bookkeeping with tracking the kernel matrix to an initial Cholesky factorization, which, in our implementation, is absorbed into the parametrization. Even though this still entails factorizations of ill-conditioned matrices, and direct solvers are also applied downstream, this empirically seems to mitigate some of the effects of ill-conditioned Jacobian matrices in solving regularized least-squares problems.
@@ Appendix B Computational tools
We perform all of our computations using Jax [10], and solve ODEs using the Tsit5 [92] solver with a PID controller for adaptive timestepping [34, 85] as recommended by, and implemented in Diffrax [47, 48]. Comparisons with SINDy use the PySINDy package [45]. Comparisons with ODR-BINDy were made using the implementation  [^3] https://github.com/llfung/ODR-BINDy from the original work [28]. Our experiments and current implementation are available in the GitHub repository [42]; an optimized implementation of JSINDy will soon be added to PySINDy. Our computations with JAX used an RTX-5000 Ada Generation GPU in a Lambda machine.